Calculate the molality for a solution prepared by mixing 125.0 mL diethyl ether (C2H6O, d = 0.736 g/cm3) with 200.0 mL acetone (C3H6O, d = 0.792 g/cm3)
1
Expert's answer
2019-11-11T07:48:02-0500

Solution.

m(dietyl ether) = 92 g = 0.092 kg n(dietyl ether) = 92/46 = 2 moles

m(acetone) = 158.4 g = 0.1584 kg n(acetone) = 158.4/58 = 2.73 moles

molarity = 4.73/0.25 = 18.92 mole/kg

Anwer. 18.92